# Restructuring Tractable Probabilistic Circuits

### v1

@@ Abstract
Probabilistic circuits (PCs) is a unifying representation for probabilistic models that support tractable inference. Numerous applications of PCs like controllable text generation depend on the ability to efficiently multiply two circuits. Existing multiplication algorithms require that the circuits respect the same structure, i.e. variable scopes decomposes according to the same vtree. In this work, we propose and study the task of restructuring structured(-decomposable) PCs, that is, transforming a structured PC such that it conforms to a target vtree. We propose a generic approach for this problem and show that it leads to novel polynomial-time algorithms for multiplying circuits respecting different vtrees, as well as a practical depth-reduction algorithm that preserves structured decomposibility. Our work opens up new avenues for tractable PC inference, suggesting the possibility of training with less restrictive PC structures while enabling efficient inference by changing their structures at inference time.
@@ 1 INTRODUCTION
A key challenge in deep generative modeling is the intractability of probabilistic reasoning (Roth,, 1996; Geh et al.,, 2024). To address this challenge, probabilistic circuits (PCs) (Darwiche,, 2003; Poon and Domingos,, 2011; Choi et al.,, 2020) has emerged as a unifying representation of tractable generative models, which support efficient and exact evaluation of various inference queries like marginalization. The tractability of PCs has now proven crucial in a range of applications, such as causal inference (Zečević et al.,, 2021; Wang and Kwiatkowska,, 2023; Busch et al.,, 2024), knowledge graph learning (Loconte et al.,, 2023) and ensuring fairness in decision making (Choi et al.,, 2021).
@@ 2 PROBABILISTIC CIRCUITS
We will use uppercase to denote variables (e.g. $X$) and lowercase to denote values of those variables (e.g. $x$). We use boldface to denote sets of variables/values (e.g. $\mathbf{X},\mathbf{x}$).
@@
A probabilistic circuit (PC) $\mathscr{A}=(\mathcal{G},\mathbf{w})$ represents a joint probability distribution over random variables $\bm{X}$ through a rooted directed acyclic (computation) graph (DAG), consisting of sum ($\oplus$), product ($\otimes$), and leaf nodes ($L$), parameterized by $\mathbf{w}$. Each node $t$ represents a probability distribution $p_{t}(\bm{X})$, defined recursively by:
@@
where $f_{t}(\bm{x})$ is a univariate input distribution function (e.g. Gaussian, Categorical), we use $\text{ch}(t)$ to denote the set of children of a node $t$, and $w_{t,c}$ is the non-negative weight associated with the edge $(t,c)$ in the DAG, which satisfy the constraint that $\sum_{c\in\text{ch}(t)}w_{t,c}=1$ for every sum node $t$. We define the scope of a node $t$ to be the variables it depends on. The function represented by a PC, denoted $p_{\mathscr{A}}(\mathbf{x})$, is the function represented by its root node; and the size of a PC, denoted $|\mathscr{A}|$, is the number of edges in its graph.
@@
A vtree $V$ over variables $\mathbf{X}$ is a rooted binary tree, where each $X\in\bm{X}$ is associated with a unique leaf node $v$ (we write $X_{v}$ for the variable associated with node $v$). Each inner node $v$ covers a set of variables $\mathbf{X}_{v}$, satisfying $\mathbf{X}_{v}=\mathbf{X}_{l}\cup\mathbf{X}_{r}$ where $l,r$ are the children of $v$. We write $V_{v}$ to denote the subtree rooted at $v$.
@@ 3 PC RESTRUCTURING
In this section, we describe a generic approach that restructures any structured-decomposable PC respecting a target vtree. The approach consists of three steps: (1) construct a Bayesian network representation of the PC; (2) find sets of latent variables in the Bayesian network that induce conditional independecies required by the target vtree; (3) construct a new structured PC recursively leveraging the conditional independence derived in (2).
@@ 3.1 Structured PCs as Bayesian Networks
Let $\mathscr{A}$ be a structured PC over variables $\mathbf{X}$ respecting vtree $V$. Given a vtree node $v\in V$, we write $\text{prod}(v)$ to denote the set of all product nodes with scope $\mathbf{X}_{v}$. We define the hidden state size $h$ of the circuit to be $\max_{v\in V}|\text{prod}(v)|$. Writing $n$ for the number of variables, the size of the circuit is then $O(nh^{2})$. [^1] The number of active sum nodes per vtree node is at most $h$, as each such node must have a different product node parent corresponding to the parent vtree node scope. This leads to $O(h^{2})$ edges per vtree node.
@@
Given a structured-decomposable and smooth PC $\mathscr{A}$ over variables $\mathbf{X}$ respecting vtree $V$, we define the augmented PC $\mathscr{A}_{\text{aug}}$ to be a copy of $\mathscr{A}$ where for each vtree node $v\in V$, we add an additional child $t_{\text{aug}}$ to each product node $t\in\text{prod}(v)$ that is a leaf node with scope $Z_{v}$ and leaf function $f_{t_{\text{aug}}}(Z_{v})=\mathds{1}_{Z_{v}=\text{idx}(t)}$.
It is not hard to see that the augmented PC $\mathscr{A}_{\text{aug}}$ is a PC over variables $\mathbf{X},\mathbf{Z}$ and retains structured decomposability and smoothness. Further, the standard marginalization algorithm for PCs ensures that the augmented PC has the correct distribution:
@@
$p_{\mathscr{A}}(\mathbf{X})=\sum_{\mathbf{z}}p_{\mathscr{A}_{\text{aug}}}(\mathbf{X},\mathbf{z})$
Let $V_{v\to Z_{v}}$ be the rooted DAG obtained by replacing all inner nodes $v$ in vtree $V$ with variable $Z_{v}$ (cf. Fig. 1). Now, we claim that the augmented PC can be interpreted as a Bayesian network with graph structure $V_{v\to Z_{v}}$. To do this, we construct a distribution $p^{*}(\mathbf{X},\mathbf{Z})$, based on the augmented PC, that factorizes as required by the Bayesian network structure.
@@
It remains to show that this distribution faithfully represents the distribution of the augmented PC, i.e. $p_{\mathscr{A}_{\text{aug}}}=p^{*}$. The intuitive idea is that each value of $\mathbf{Z}$ corresponds to a subtree of $\mathscr{A}_{\text{aug}}$, whose value is precisely given by the product of weights and leaf functions specified by the Bayesian network; we refer readers to the Appendix for the complete proof. We thus have the following mapping from structured PCs to tree-shaped Bayesian networks:
@@
Let $\mathscr{A}$ be a structured-decomposable and smooth PC over variables $\mathbf{X}$ respecting vtree $V$. Then there exists a Bayesian network $G_{\mathscr{A}}$ over variables $\mathbf{X}$ and $\mathbf{Z}=\{Z_{v}|v\in V\}$ with graph $V_{v\mapsto Z_{v}}$ such that $\sum_{z}p_{G}(\mathbf{X},\bm{z})=p_{\mathscr{A}}(\mathbf{X})$.
Since we have shown that $p_{\mathscr{A}}$ and $p_{G}$ represents the same distribution over the observed variables $\mathbf{X}$, we will drop the subscripts when there is no ambiguity.
@@ 3.2 Recursive PC Restructuring
Suppose we have a PC $\mathscr{A}$ with its Bayesian network representation $G_{\mathscr{A}}$ and vtree $V$, and let $W$ be some other vtree. We now show how to construct a new PC respecting $W$ that encodes the same distribution as $\mathscr{A}$. The rough idea is to label each vtree node $w\!\in\!W$ with a subset of latent variables $\mathbf{C}_{w}\!\subseteq\!G_{\mathscr{A}}$ such that $\mathbf{X}_{w}$ is conditionally independent from $\mathbf{X}\setminus\mathbf{X}_{w}$ given $\mathbf{C}_{w}$. To characterize such properties, we introduce covers:
@@
Given a tree-shaped Bayesian network $G_{\mathscr{A}}$ as constructed in Sec. 3.1, we say that $\mathbf{C}\subseteq\mathbf{Z}$ covers $\mathbf{S}\subseteq\mathbf{X}$ if $\mathbf{C}$ blocks222a path $P$ is blocked by a set $\mathbf{S}$ if $P\cap\mathbf{S}\neq\varnothing$. all paths between $\mathbf{S}$ and $\mathbf{X}\!\setminus\!\mathbf{S}$ in $G_{\mathscr{A}}$.
@@
$\mathbf{A},\mathbf{B}\!\subseteq\!G_{\mathscr{A}}$ are conditionally independent given $\mathbf{C}\!\subseteq\!G_{\mathscr{A}}$ if and only if $\mathbf{C}$ blocks all paths between $\mathbf{A}$ and $\mathbf{B}$. In particular, if $\mathbf{C}$ covers $\mathbf{S}$ then $\mathbf{S}$ and $\mathbf{X}\!\setminus\!\mathbf{S}$ are conditionally independent given $\mathbf{C}$.
Our goal is to recursively construct vectors of sum nodes $\oplus_{i}$ representing the probability distributions $p(\mathbf{X}_{w}\>|\>\mathbf{C}_{w}\!=\!i)$. Letting $l$ and $r$ be the children of $w$, we will establish a recurrence relation between $p(\mathbf{X}_{w}\>|\>\mathbf{C}_{w})$, $p(\mathbf{X}_{l}\>|\>\mathbf{C}_{l})$ and $p(\mathbf{X}_{r}\>|\>\mathbf{C}_{r})$. This requires the vtree labels to satisfy the following properties:
@@
Given the Bayesian network $G_{\mathscr{A}}$ and target vtree $W$, a valid labelling of $W$ with respect to $G_{\mathscr{A}}$ associates each node $w\!\in\!W$ with a subset of latent variables $\mathbf{C}_{w}\!\subseteq\!G_{V}$ s.t.
@@
    $\mathbf{C}_{w}$ covers $\mathbf{X}_{w}$ in $G_{\mathscr{A}}$.
@@
    $\mathbf{C}_{l}$ blocks all paths between $\mathbf{X}_{l}$ and $\mathbf{C}_{r}\cup\mathbf{C}_{w}$.
@@
    $\mathbf{C}_{r}$ blocks all paths between $\mathbf{X}_{r}$ and $\mathbf{C}_{l}\cup\mathbf{C}_{w}$.
Furthermore, w.l.o.g., we set $\mathbf{C}_{\text{root of }W}\!:=\!\varnothing$ and $\mathbf{C}_{X_{j}}\!:=\!\text{parent of }X_{j}\text{ in }G_{\mathscr{A}}$ for the leaf nodes $X_{j}\!\in\!W$. See Figure 1(c) for an example.
Figure 2: Recursive construction of vectors of sum nodes representing $p(\mathbf{X}_{w}\>|\>\mathbf{C}_{w})$
Assuming that we have computed a valid labelling for $W$, we can then proceed to construct the desired PC by a bottom-up recursion on $W$. For the base case, if $w$ is a leaf node representing some random variable $X_{j}$, $p(X_{j}\>|\>\mathbf{C}_{X_{j}})=p(X_{j}\>|\>\text{parent of }X_{j}\text{ in }G_{\mathscr{A}})$, which is directly given by the conditional probability table of $G_{\mathscr{A}}$. For the induction step, when $w$ is a inner node with children $l$ and $r$, we have the recurrence relation:
$p(\mathbf{X}_{w}\>|\>\mathbf{C}_{w})$
$=\sum_{(\mathbf{C}_{l}\cup\mathbf{C}_{r})\setminus\mathbf{C}_{w}}p(\mathbf{X}_{l},\mathbf{X}_{r}\>|\>\mathbf{C}_{l},\mathbf{C}_{r})\cdot p(\mathbf{C}_{l},\mathbf{C}_{r}\>|\>\mathbf{C}_{w})$
$=\sum_{(\mathbf{C}_{l}\cup\mathbf{C}_{r})\setminus\mathbf{C}_{w}}p(\mathbf{X}_{l}\>|\>\mathbf{C}_{l})\cdot p(\mathbf{X}_{r}\>|\>\mathbf{C}_{r})\cdot p(\mathbf{C}_{l},\mathbf{C}_{r}\>|\>\mathbf{C}_{w})$
Here the first step follows from Property 2 and 3, and the second step follows from all properties in Defintion 3.6. The circuit materialization of the recurrence relation is shown in Figure 2. Note that if $w$ is the root, then $p(\mathbf{X}_{w}\>|\>\mathbf{C}_{w})$ becomes $p(\mathbf{X})$, which is a single sum node representing the distribution of $\mathscr{A}$. The complete recursion is given by Algorithm 1.
@@
     $\bigoplus_{\mathbf{X}_{l},\mathbf{C}_{l}}\!\leftarrow\!\textsc{ConstructCircuit}(l)$
     $\bigoplus_{\mathbf{X}_{r},\mathbf{C}_{r}}\!\leftarrow\!\textsc{ConstructCircuit}(r)$
     $\bigoplus_{\mathbf{X}_{w},\mathbf{C}_{w}}\!\leftarrow\!\sum\limits_{(\mathbf{C}_{l}\cup\mathbf{C}_{r})\setminus\mathbf{C_{w}}}\!\bigoplus_{\mathbf{C}_{l}}\cdot\bigoplus_{\mathbf{C}_{r}}\cdot p(\mathbf{C}_{l},\mathbf{C}_{r}\>|\>\mathbf{C}_{w})$
     return $\bigoplus_{\mathbf{C}_{w}}$
@@
Let $h$ be the number of hidden states of the original PC $\mathscr{A}$ and $n$ the number of random variables. The number of hidden states of the restructured PC is given by $O(h^{M})$ where $M\!=\!{\max}_{w\in W}|\mathbf{C}_{l}\cup\mathbf{C}_{r}|$ and the size of the restructured PC is bounded by $O(nh^{M^{\prime}})$ where $M^{\prime}\!=\!\max_{w\in W}|\mathbf{C}_{l}\cup\mathbf{C}_{r}\cup\mathbf{C}_{w}|\leq 2M$. We refer to $M^{\prime}$ as the cardinality of the labelling $\mathbf{C}_{w}$.
@@
Let $\mathscr{A}^{\prime}$ be the restructured circuit respecting $W$. As described in Algorithm 1, for each inner node $w\in W$, we construct two layers of nodes as shown in Figure 2. By construction, the product layer contains all product nodes respecting the vtree node $w$ and its cardinality is given by $O(h^{|\mathbf{C}_{l}\cup\mathbf{C}_{r}|})$; we set $M\!:=\!\max_{w\in W}|\mathbf{C}_{l}\cup\mathbf{C}_{r}|$ and it follows that the hidden states size of $\mathscr{B}$ is given by $O(h^{M})$. Similarly, the number of edges in the sum layer is given by $O(h^{|\mathbf{C}_{l}\cup\mathbf{C}_{r}\cup\mathbf{C}_{w}|})$ and the number of product edges is given by $O(h^{|\mathbf{C}_{l}\cup\mathbf{C}_{r}|})$; since there are $O(n)$ vtree nodes in total, the total number of edges in $\mathscr{B}$ is given by $O(nh^{M^{\prime}})$, with $M^{\prime}=\max_{w\in W}|\mathbf{C}_{l}\cup\mathbf{C}_{r}\cup\mathbf{C}_{w}|$. ∎

### v2

@@ Abstract
Probabilistic circuits (PCs) are a unifying representation for probabilistic models that support tractable inference. Numerous applications of PCs like controllable text generation depend on the ability to efficiently multiply two circuits. Existing multiplication algorithms require that the circuits respect the same structure, i.e. variable scopes decomposes according to the same vtree. In this work, we propose and study the task of restructuring structured(-decomposable) PCs, that is, transforming a structured PC such that it conforms to a target vtree. We propose a generic approach for this problem and show that it leads to novel polynomial-time algorithms for multiplying circuits respecting different vtrees, as well as a practical depth-reduction algorithm that preserves structured decomposibility. Our work opens up new avenues for tractable PC inference, suggesting the possibility of training with less restrictive PC structures while enabling efficient inference by changing their structures at inference time.
@@ 1 INTRODUCTION
A key challenge in deep generative modeling is the intractability of probabilistic reasoning (Roth,, 1996; Geh et al.,, 2024). To address this challenge, probabilistic circuits (PCs) (Darwiche,, 2003; Poon and Domingos,, 2011; Choi et al.,, 2020) has emerged as a unifying representation of tractable generative models for high-dimensional probability distributions. In particular, PCs support efficient and exact evaluation of various inference queries such as marginalization. The tractability of PCs has proven crucial in various applications, such as causal inference (Zečević et al.,, 2021; Wang and Kwiatkowska,, 2023; Busch et al.,, 2024), knowledge graph learning (Loconte et al.,, 2023) and ensuring fairness in decision making (Choi et al.,, 2021).
@@
In summary, our key contributions are: (1) we identify the restructuring problem, which concerns the computational complexity of converting between different tractable PC structures; (2) we present a general algorithm for mapping between any two structures (Section 3); (3) we show that restructuring is possible in (quasi-)polynomial complexity for a class of contiguous structures (Section 4); and (4) we illustrate the application of restructuring to two important problems: multiplying probability distributions (Section 4) and reducing the depth of the model structure (Section 5), where we derive novel results on tractability.
@@ 2 PROBABILISTIC CIRCUITS
We will use uppercase to denote variables (e.g. $X$) and lowercase to denote values of those variables (e.g. $x$). We use boldface to denote sets of variables/values (e.g. $\bm{X},\bm{x}$).
@@
A probabilistic circuit (PC) $\mathscr{A}=(\mathcal{G},\bm{w})$ represents a joint probability distribution over random variables $\bm{X}$ through a rooted directed acyclic (computation) graph (DAG), consisting of sum ($\oplus$), product ($\otimes$), and leaf nodes ($L$), parameterized by $\bm{w}$. Each node $t$ represents a probability distribution $p_{t}(\bm{X})$, defined recursively by:
@@
where $f_{t}(\bm{x})$ is a univariate input distribution function (e.g. Gaussian, Categorical), we use $\text{ch}(t)$ to denote the set of children of a node $t$, and $w_{t,c}$ is the non-negative weight associated with the edge $(t,c)$ in the DAG, which satisfy the constraint that $\sum_{c\in\text{ch}(t)}w_{t,c}=1$ for every sum node $t$. We define the scope of a node $t$ to be the variables it depends on. The function represented by a PC, denoted $p_{\mathscr{A}}(\bm{x})$, is the function represented by its root node; and the size of a PC, denoted $|\mathscr{A}|$, is the number of edges in its graph.
@@
A vtree $V$ over variables $\bm{X}$ is a rooted binary tree, where each $X\in\bm{X}$ is associated with a unique leaf node $v$ (we write $X_{v}$ for the variable associated with node $v$). Each inner node $v$ covers a set of variables $\bm{X}_{v}$, satisfying $\bm{X}_{v}=\bm{X}_{l}\cup\bm{X}_{r}$ where $l,r$ are the children of $v$. We write $V_{v}$ to denote the subtree rooted at $v$.
@@ 3 PC RESTRUCTURING
In this section, we describe a generic approach that restructures any structured-decomposable PC to respect a target vtree. The approach consists of three steps: (1) construct a Bayesian network representation of the PC; (2) find sets of latent variables in the Bayesian network that induce conditional independecies required by the target vtree; (3) construct a new structured PC recursively leveraging the conditional independence derived in (2).
@@ 3.1 Structured PCs as Bayesian Networks
Let $\mathscr{A}$ be a structured PC over variables $\bm{X}$ respecting vtree $V$. Given a vtree node $v\in V$, we write $\text{prod}(v)$ to denote the set of all product nodes with scope $\bm{X}_{v}$. We define the hidden state size $h$ of the circuit to be $\max_{v\in V}|\text{prod}(v)|$. Writing $n$ for the number of variables, the size of the circuit is then $O(nh^{2})$. [^1] The number of active sum nodes per vtree node is at most $h$, as each such node must have a different product node parent corresponding to the parent vtree node scope. This leads to $O(h^{2})$ edges per vtree node.
@@
Given a structured-decomposable and smooth PC $\mathscr{A}$ over variables $\bm{X}$ respecting vtree $V$, we define the augmented PC $\mathscr{A}_{\text{aug}}$ to be a copy of $\mathscr{A}$ where for each vtree node $v\in V$, we add an additional child $t_{\text{aug}}$ to each product node $t\in\text{prod}(v)$ that is a leaf node with scope $Z_{v}$ and leaf function $f_{t_{\text{aug}}}(Z_{v})=\mathds{1}_{Z_{v}=\text{idx}(t)}$.
It is not hard to see that the augmented PC $\mathscr{A}_{\text{aug}}$ is a PC over variables $\bm{X},\bm{Z}$ and retains structured decomposability and smoothness. Further, the standard marginalization algorithm for PCs ensures that the augmented PC has the correct distribution:
@@
$p_{\mathscr{A}}(\bm{X})=\sum_{\bm{z}}p_{\mathscr{A}_{\text{aug}}}(\bm{X},\bm{z})$
Let $V_{v\to Z_{v}}$ be the rooted DAG obtained by replacing all inner nodes $v$ in vtree $V$ with variable $Z_{v}$ (cf. Fig. 1). Now, we claim that the augmented PC can be interpreted as a Bayesian network with graph structure $V_{v\to Z_{v}}$. To do this, we construct a distribution $p^{*}(\bm{X},\bm{Z})$, based on the augmented PC, that factorizes as required by the Bayesian network structure.
@@
It remains to show that this distribution faithfully represents the distribution of the augmented PC, i.e. $p_{\mathscr{A}_{\text{aug}}}=p^{*}$. The intuitive idea is that each value of $\bm{Z}$ corresponds to a subtree of $\mathscr{A}_{\text{aug}}$, whose value is precisely given by the product of weights and leaf functions specified by the Bayesian network; we refer readers to the Appendix for the complete proof. We thus have the following mapping from structured PCs to tree-shaped Bayesian networks:
@@
Let $\mathscr{A}$ be a structured-decomposable and smooth PC over variables $\bm{X}$ respecting vtree $V$. Then there exists a Bayesian network $G_{\mathscr{A}}$ over variables $\bm{X}$ and $\bm{Z}=\{Z_{v}|v\in V\}$ with graph $V_{v\mapsto Z_{v}}$ such that $\sum_{z}p_{G}(\bm{X},\bm{z})=p_{\mathscr{A}}(\bm{X})$.
Since we have shown that $p_{\mathscr{A}}$ and $p_{G}$ represents the same distribution over the observed variables $\bm{X}$, we will drop the subscripts when there is no ambiguity.
@@ 3.2 Recursive PC Restructuring
Suppose we have a PC $\mathscr{A}$ with its Bayesian network representation $G_{\mathscr{A}}$ and vtree $V$, and let $W$ be some other vtree. We now show how to construct a new PC respecting $W$ that encodes the same distribution as $\mathscr{A}$. The rough idea is to label each vtree node $w\!\in\!W$ with a subset of latent variables $\bm{C}_{w}\!\subseteq\!G_{\mathscr{A}}$ such that $\bm{X}_{w}$ is conditionally independent from $\bm{X}\setminus\bm{X}_{w}$ given $\bm{C}_{w}$. To characterize such properties, we introduce covers:
Definition 3.4 (Blocked Path).
Given a directed rooted tree, we say that a path $P$ is blocked by a set of nodes $\bm{S}$ if $P\cap\bm{S}\neq\varnothing$.
@@
Given a tree-shaped Bayesian network $G_{\mathscr{A}}$ as constructed in Sec. 3.1, we say that $\bm{C}\subseteq\bm{Z}$ covers $\bm{S}\subseteq\bm{X}$ if $\bm{C}$ blocks all paths between $\bm{S}$ and $\bm{X}\!\setminus\!\bm{S}$ in $G_{\mathscr{A}}$.
@@
$\bm{A},\bm{B}\!\subseteq\!G_{\mathscr{A}}$ are conditionally independent given $\bm{C}\!\subseteq\!G_{\mathscr{A}}$ if and only if $\bm{C}$ blocks all paths between $\bm{A}$ and $\bm{B}$ [^2] Defining blocked paths for DAGs requires considering colliders, which do not occur in directed rooted trees. In particular, if $\bm{C}$ covers $\bm{S}$ then $\bm{S}$ and $\bm{X}\!\setminus\!\bm{S}$ are conditionally independent given $\bm{C}$.
Our goal is to recursively construct vectors of sum nodes $\oplus_{i}$ representing the probability distributions $p(\bm{X}_{w}\>|\>\bm{C}_{w}\!=\!i)$. Letting $l$ and $r$ be the children of $w$, we will establish a recurrence relation between $p(\bm{X}_{w}\>|\>\bm{C}_{w})$, $p(\bm{X}_{l}\>|\>\bm{C}_{l})$ and $p(\bm{X}_{r}\>|\>\bm{C}_{r})$. This requires the vtree labels to satisfy the following properties:
@@
Given the Bayesian network $G_{\mathscr{A}}$ and target vtree $W$, a valid labelling of $W$ with respect to $G_{\mathscr{A}}$ associates each node $w\!\in\!W$ with a subset of latent variables $\bm{C}_{w}\!\subseteq\!G_{V}$ s.t.
@@
    $\bm{C}_{w}$ covers $\bm{X}_{w}$ in $G_{\mathscr{A}}$.
@@
    $\bm{C}_{l}$ blocks all paths between $\bm{X}_{l}$ and $\bm{C}_{r}\cup\bm{C}_{w}$.
@@
    $\bm{C}_{r}$ blocks all paths between $\bm{X}_{r}$ and $\bm{C}_{l}\cup\bm{C}_{w}$.
Furthermore, w.l.o.g., we set $\bm{C}_{\text{root of }W}\!:=\!\varnothing$ and $\bm{C}_{X_{j}}\!:=\!\text{parent of }X_{j}\text{ in }G_{\mathscr{A}}$ for the leaf nodes $X_{j}\!\in\!W$. See Figure 1(c) for an example.
Figure 2: Recursive construction of vectors of sum nodes representing $p(\bm{X}_{w}\>|\>\bm{C}_{w})$
Assuming that we have computed a valid labelling for $W$, we can then proceed to construct the desired PC by a bottom-up recursion on $W$. For the base case, if $w$ is a leaf node representing some random variable $X_{j}$, $p(X_{j}\>|\>\bm{C}_{X_{j}})=p(X_{j}\>|\>\text{parent of }X_{j}\text{ in }G_{\mathscr{A}})$, which is directly given by the conditional probability table of $G_{\mathscr{A}}$. For the induction step, when $w$ is a inner node with children $l$ and $r$, we have the recurrence relation:
$p(\bm{X}_{w}\>|\>\bm{C}_{w})$
$=\sum_{(\bm{C}_{l}\cup\bm{C}_{r})\setminus\bm{C}_{w}}p(\bm{X}_{l},\bm{X}_{r}\>|\>\bm{C}_{l},\bm{C}_{r})\cdot p(\bm{C}_{l},\bm{C}_{r}\>|\>\bm{C}_{w})$
$=\sum_{(\bm{C}_{l}\cup\bm{C}_{r})\setminus\bm{C}_{w}}p(\bm{X}_{l}\>|\>\bm{C}_{l})\cdot p(\bm{X}_{r}\>|\>\bm{C}_{r})\cdot p(\bm{C}_{l},\bm{C}_{r}\>|\>\bm{C}_{w})$
Here the first step follows from Property 2 and 3, and the second step follows from all properties in Defintion 3.7. The circuit materialization of the recurrence relation is shown in Figure 2. Note that if $w$ is the root, then $p(\bm{X}_{w}\>|\>\bm{C}_{w})$ becomes $p(\bm{X})$, which is a single sum node representing the distribution of $\mathscr{A}$. The complete recursion is given by Algorithm 1.
@@
     $\bigoplus_{\bm{X}_{l},\bm{C}_{l}}\!\leftarrow\!\textsc{ConstructCircuit}(l)$
     $\bigoplus_{\bm{X}_{r},\bm{C}_{r}}\!\leftarrow\!\textsc{ConstructCircuit}(r)$
     $\bigoplus_{\bm{X}_{w},\bm{C}_{w}}\!\leftarrow\!\sum\limits_{(\bm{C}_{l}\cup\bm{C}_{r})\setminus\bm{C_{w}}}\!\bigoplus_{\bm{C}_{l}}\cdot\bigoplus_{\bm{C}_{r}}\cdot p(\bm{C}_{l},\bm{C}_{r}\>|\>\bm{C}_{w})$
     return $\bigoplus_{\bm{C}_{w}}$
@@
Let $h$ be the number of hidden states of the original PC $\mathscr{A}$ and $n$ the number of random variables. The number of hidden states of the restructured PC is given by $O(h^{M})$ where $M\!=\!{\max}_{w\in W}|\bm{C}_{l}\cup\bm{C}_{r}|$ and the size of the restructured PC is bounded by $O(nh^{M^{\prime}})$ where $M^{\prime}\!=\!\max_{w\in W}|\bm{C}_{l}\cup\bm{C}_{r}\cup\bm{C}_{w}|\leq 2M$. We refer to $M^{\prime}$ as the cardinality of the labelling $\bm{C}_{w}$.
@@
Let $\mathscr{A}^{\prime}$ be the restructured circuit respecting $W$. As described in Algorithm 1, for each inner node $w\in W$, we construct two layers of nodes as shown in Figure 2. By construction, the product layer contains all product nodes respecting the vtree node $w$ and its cardinality is given by $O(h^{|\bm{C}_{l}\cup\bm{C}_{r}|})$; we set $M\!:=\!\max_{w\in W}|\bm{C}_{l}\cup\bm{C}_{r}|$ and it follows that the hidden states size of $\mathscr{B}$ is given by $O(h^{M})$. Similarly, the number of edges in the sum layer is given by $O(h^{|\bm{C}_{l}\cup\bm{C}_{r}\cup\bm{C}_{w}|})$ and the number of product edges is given by $O(h^{|\bm{C}_{l}\cup\bm{C}_{r}|})$; since there are $O(n)$ vtree nodes in total, the total number of edges in $\mathscr{B}$ is given by $O(nh^{M^{\prime}})$, with $M^{\prime}=\max_{w\in W}|\bm{C}_{l}\cup\bm{C}_{r}\cup\bm{C}_{w}|$. ∎